Brief Exercise 18-06

Determine the missing amounts.

Unit Selling
Price

Unit Variable
Costs

Unit Contribution
Margin

Contribution Margin
Ratio

1. $800 $416 $enter a dollar amount

(a) enter percentages

% (b)
2. $250 $enter a dollar amount

(c) $85 enter percentages

% (d)
3. $enter a dollar amount

(e) $enter a dollar amount

(f) $455 35 %
Answer #1

Answer

--Requirement

Unit selling price Unit Variable cost Unit Contribution margin CM Ratio
1 $800 $416 $384 0.48
2 $250 $165 $85 0.34
3 $1,300 $845 $455 35%

--Working

Unit selling price Unit Variable cost Unit Contribution margin CM Ratio
1 800 416 =800-416 =384/800
2 250 =250-85 85 =85/250
3 =455/(0.35) =1300-455 455 0.35